OpenStudy (anonymous):

Prove the area of a parallelogram is 1/2 d_1 d_2 sin(theta).

OpenStudy (anonymous):

Is there a figure? If d_1 and d_2 are the side lengths, this isn't quite right

OpenStudy (anonymous):

They are the diagonals.

OpenStudy (anonymous):

Can you draw a figure, or shall I?

OpenStudy (anonymous):

And what is theta?

OpenStudy (anonymous):

E

OpenStudy (anonymous):

While you're thinking, I'll point out that the figure already has the diagonal segments labeled to show that at the center, each diagonal is divided in half. You can prove that easily from the fact that the sides are parallel. So the area of triangle DAC is 1/2 the total.

OpenStudy (anonymous):

You've got me stumped for the moment, better ask for help.

OpenStudy (anonymous):

My main issue is: what does sin(theta) give me, the height? The base?

OpenStudy (anonymous):

Right. I think of rearranging the pieces, and get lost. I think of direct calculation and it gets to be a mess. I think of the law of sines, and hope for cancellations. That'd be my bet.

OpenStudy (anonymous):

How about this Let d_1 = the diagonal with a single mark, and d_2 the other one. Let b = the base of the parallelogram (DC) and s the side (BC). Let theta = the angle EDC. Now sin theta / (1/2) d_2 = sin E / b

OpenStudy (anonymous):

And the area of BDC = d_1 sin theta times b

OpenStudy (anonymous):

Of a triangle, the area of .5 a b sin(c). .5 of diagonal one and 2 * sin (E) gives me one of the triangle areas. sin(180-E)=sin(E), so we can find the area of the other triangle, and multiply the sum of those two triangles by two.

OpenStudy (anonymous):

... to get the total area.

OpenStudy (anonymous):

It's sounding good.

OpenStudy (anonymous):

I'm having a problem of a factor of 1/2. Doing it my way: sin theta / (d_2/2) = sin E / b sin theta = (d_2/2) (1/b) sin E Area of DBC (half the whole) = d_1 b sin theta A (1/d_1) (1/b) = sin theta = (d_2/2) (1/b) sin E cancel (1/b) A = d_1 (d_2/2) sin E But that's only half the area. total = d_1 d_2 sin E What do you think?

OpenStudy (anonymous):

Wanted to show another way. I cut the solid into 4 pieces and rearranged them. See the figure. Standard result with vectors is that the area is A x B which here is d_1 d_2/2 sin E http://en.wikipedia.org/wiki/Cross_product

OpenStudy (anonymous):

Use the second figure, first one's bad. Was fun!